OpenStudy (anonymous):

FIND DERIVATE (3T)(SINT / T) - ( T^2 / 3)

OpenStudy (anonymous):

wow you have product rule, and quotient. Ive never seen that before but its pretty simple

OpenStudy (anonymous):

\[(3T)(SINT / T) - ( T^2 / 3) \] it is that your equation?

OpenStudy (anonymous):

YES I KNOW IT SEEMS EASY BUT THAT COULD NOT BE AN EASY ANSWER MY PROFFESOR LIKES TO GIVE TRICKY QUESTIONS

OpenStudy (anonymous):

3cost - 2/3(t)

OpenStudy (anonymous):

(3T)[(T)(cos(t)-(1)(sin(t)]/t^2]+(3)(sint/t))-(3)(2t)-(t^2)(0)/9

OpenStudy (anonymous):

its such a crazy combination but that should be right

OpenStudy (anonymous):

I GET 3COST BUT NOT -2/3T

OpenStudy (anonymous):

isnt sin t/t=1?

OpenStudy (anonymous):

the derivative of 3sint = 3cost derivative of (-t^2/3) = -2/3 (t) together 3cost-(2/3)t

OpenStudy (anonymous):

(3T)(1)-(T^/3) = 3-[3(1)/9]=3-1/3=3/3-1/3=2/3

OpenStudy (anonymous):

the answer is just 2/3 rewrite it

OpenStudy (anonymous):

SINT/T is an identity which can be rewritten as =1

OpenStudy (anonymous):

we are talking about the derivative... not limits

OpenStudy (anonymous):

for limits we know that sint/t = 1 but the question said FIND THE DERIVATIVE

OpenStudy (anonymous):

yea then you are right 3cosT-2/3T

OpenStudy (anonymous):

OK I THINK THIS IS WERE I WAS CONFUSED IS NOT -T^ RAISED TO THE 2/3 I KNOW U BRING THE 2/3 BEFORE THE T

OpenStudy (anonymous):

IS T^2 THIS TERM DIVIDED BY 3

OpenStudy (anonymous):

shouldnt it still be raised to the -1/3

OpenStudy (anonymous):

TO FIND DERIVATE SO DONT WE HAVE TO MULTIPLY KINDA OF 3T THE OUTSIDE (SINT/T) MINUS (T^2 DIVIDED BY 3)